OpenStudy (anonymous):

A rectangle has a perimeter of 26 feet. Twice its length (l) is one foot more than three times its width (w). What is the length of the rectangle, in feet? A. 5 B. 13 C. 8 D. 16

OpenStudy (anonymous):

perimeter of rectangle = 2*(Length+width) as per problem 2L+1 = 3W where L = length and W = width plug the value of width in the formula for perimeter hope u can determine length now

OpenStudy (texaschic101):

P = 2L + 2W P = 26 2L = 3W + 1...... L = 3/2W + 1/2 now sub 26 = 2(3/2W + 1/2) + 2W can you solve it from here ? First solve for W, then sub that back in and solve for L.